# EVA-Score: Evaluation of Long-form Summarization on Informativeness through Extraction and Validation

### v1

@@ 5.1 Dataset
To test the applicability of EVA-Score, we collect long-form summarization with different context lengths. Specifically, we adopt CNN Dailymail(See et al., 2017), PubMed, GovReport(Cohan et al., 2018), arXiv, BookSum(Kryściński et al., 2022). In particular, we use the chapter-level summarization version of BookSum instead of the book level for less difficulty. For all these datasets, we use validation datasets for evaluation.
Dataset CNN Dailymail Pubmed arXiv Gov Report BookSum
Average Text Length 781 3111 5894 7797 14337
Average Summary Length 56 216 172 573 570
Table 1: The Average Text Length and the Average Summary Length in our experimented dataset.
5.2 Model
To perform a fair comparison, we experiment on a wide range of Generative Language Models of different numbers of parameters, and from different architectures. For open-source models, we test EVA-Score on Mistral-7B(Jiang et al., 2023) which is a MoE model, variants of Llama-2-chat(Touvron et al., 2023), with parameters 7B and 13B, variants of Vicuna(Zheng et al., 2023) with parameters 7B and 13B and Gemma-7B-it(Team et al., 2024). For close-source models, we experiment on ChatGPT and GPT-4(0125-preview version).
@@ 5.3 Baseline
We compare the results of EVA-Score with traditional evaluation metrics like ROUGE and BERTScore(Zhang et al., 2020) which are based on text overlap and word embeddings. Recently, many works have leveraged LLMs to obtain the informativeness of a given text corresponding to a reference. We include Vicuna-13B and Mixtral-7B-Instruct as our open-source LLM evaluators and ChatGPT and GPT-4(0125-preview version) as our close-source LLM evaluators. For all the LLMs, we instruct them using the prompt from Table 12, imitating the human annotation process.
@@ 5.4 Human Annotation
We are committed to developing an evaluation metric that resonates with human judgment. To achieve this, we randomly select 30 records from the dataset and assess the informativeness of each record by ourselves. Initially, we identify sentence-level and document-level atomic facts, denoted as $S$ and $\hat{S}$, from $s$ and $\hat{s}$, respectively. They then determine the overlap between $S$ and $\hat{S}$. Redundant information, which appears multiple times, is recorded only once. Subsequently, we calculate Precision and Recall for these overlaps and derive the F1 Score as their harmonic mean. Furthermore, we ensure that each record is evaluated by at least two annotators to maintain assessment fairness.
@@ 5.5 Main Results
Model-based Evaluation Metric
Mixtral-7B-Instruct 0.097 0.052 0.030 -0.653 -0.400 -0.333
Vicuna-13B 0.065 -0.040 -0.029 0.367 0 0
ChatGPT 0.138 0.210 0.162 0.430 0.600 0.333
@@
EVA-Score (Ours) 0.710 0.673 0.503 0.975 1.000 1.000
Table 2: Text-level and system-level Pearson($\sigma$), Spearman($\gamma$), and Kendall($\tau$) correlations with different evaluation metrics. The highest correlation among the methods is bold
In this section, we report the correlation between each evaluation metric and human annotations. We adopt text-level and system-level Pearson($\sigma$), Spearman($\gamma$), and Kendall($\tau$) correlation. Text-level correlation is calculated by computing the correlation coefficients between human evaluations and automatic evaluation metrics and then averaging these coefficients for all generated texts per instruction. System-level correlation, on the other hand, is determined by computing the correlation coefficients between human evaluations and automatic evaluation metrics, but this time for the average score of each dataset.
From Table 2, we observe that our EVA-Score outperforms all other model-based and traditional evaluation metrics in achieving the highest correlation with human judgments across all indicators at both levels. At the System Level, EVA-Score achieves perfect Spearman and Kendall correlations of 1.00, demonstrating complete alignment with human assessments. In contrast, Traditional Evaluation Methods, while somewhat correlated with human evaluations, generally lack strong correlations, highlighting their limitations.
Surprisingly, we find that the information scores assigned by Vicuna-13B show little correlation with humans, and At the System Level, the scores from Mistral-7B show even negative correlation with humans. We attribute this to the situation that the informativeness is obtained by quantitative calculation using EVA-Score and Human Annotation. The instruction may be too complex for them to deal with within a single interaction which requires them first to extract atomic facts and document-level relations and verify all of them one by one. However, for models like ChatGPT and GPT-4, while their correlation with human judgments at the Text Level remains minimal, they exhibit promising outcomes at the System Level using only In-Context Learning.
We also report our results on several long-form summarization datasets, including CNN Dailymail, PubMed, GovReport, arXiv, and BookSum, using three different evaluation metrics, i.e. EVA-Score, ROUGE-1, and BERTScore. Since this work focuses on the evaluation of long-form summarization, we don’t pursue feeding long context into LLMs without information loss. Instead, we believe that useful information mostly appears at the beginning or the end of a text, we just simply truncate the context from both the beginning and end to meet the models’ max input sequence length.
@@
Table 3: The results of Several LLMs on different long-form summarization datasets. We denote EVA-Score as EVA, R-1 as ROUGE-1, BS as BERTScore. All the values are normalized between 0 and 1. We indicate the highest score of EVA-Score using bold.
Referring to Table 3, we observe that as the context lengthens, all tested models experience varying degrees of performance decline. Particularly notable is the performance at the part-of-book level, i.e. chapter summarization in BookSum, where the results are markedly poor compared to the ground truth. This issue, however, can be mitigated by employing methods that feed the entire context into the model, rather than truncating it as we have done in our initial experiments. Additionally, compared to other evaluation metrics, the EVA-Score demonstrates a more pronounced and consistent decline in performance as context length increases, with fewer fluctuations. This pattern aligns well with human cognitive processes, indicating that as information load increases, the challenge of maintaining high summarization quality also grows. This consistent pattern of decline not only reflects realistic expectations of model performance but also underscores the effectiveness of the EVA-Score in capturing these dynamics, further validating its utility from another perspective.
Moreover, models using ROPE(Su et al., 2023) or other methods, enabling longer input sequences show greater performance in long-form summarization when the context length becomes larger, mainly due to their ability to comprehend a long context or search in a long context. However, when the context is not so long, they show comparable or inferior performance with assistant models with normal sequence length, e.g. 4096 or 8192.
Under the evaluation framework of the EVA-Score, GPT-4 demonstrates robust capabilities, particularly with longer input sequences, compared to other models. In the CNN Dailymail dataset, it trails behind models such as Vicuna-7B and ChatGPT. This is primarily due to GPT-4’s tendency to generate more extensive summaries than its counterparts in settings with relatively short contexts, thereby increasing its count of atomic facts. While this might allow it to cover more points than other models, it often results in lower precision due to the inclusion of information that extends beyond the reference summary. Nevertheless, in other datasets such as arXiv and Gov Report, GPT-4 outperforms all competitors. It also secures a second-place ranking in Pubmed and BookSum, affirming its effectiveness across diverse text summarization tasks.
To be mentioned, Mixtral-7B-Instruct, compared with other models with the same parameter or even larger, shows great potential to fulfill long-form summarization and even beats GPT-4 by a small margin in chapter-level BookSum, which meets the findings of Chang et al. (2024).
@@ 5.6 Analysis of Summarization Length
Indicator Avg Length Avg Length Avg Length Avg Length Avg Length
@@
Llama-2-7B-chat-hf 118 143 108 140 185
Llama-2-13B-chat-hf 124 184 221 236 179
@@
Table 4: The average summarization length(denoted by Avg Length by different model. The longest is marked underline
To elucidate why GPT-4 underperforms on the EVA-Score with a shorter context, we conduct an analysis focused on the summarization lengths produced by different models, as detailed in Table 4. Models pre-trained with longer input sequences tend to generate more extensive responses. Specifically, GPT-4 produces summaries of 153 tokens in CNN Dailymail, 227 tokens in Pubmed, and 272 tokens in arXiv and Mixtral-7B-Instruct outputs 353 tokens in Gov Report and 251 tokens in BookSum whose maximum input sequences length are 128000 and 32768 respectively.
Upon manual examination, we observe that shorter contexts typically result in shorter ground truth. Consequently, if a model generates a lengthy prediction, this verbosity is penalized with a reduced EVA-Score due to diminished precision. As the context lengthens, references become richer in information. However, these models tend to maintain a similar response length, lacking significant information addition, thereby failing to capture enough relevant points, especially in longer predictions. This discrepancy explains GPT-4’s shortfall in CNN Dailymail and the superior performance of Mistral-7B-Instruct, despite having a comparable number of parameters.
Furthermore, GPT-4, which is refined through Reinforcement Learning from Human Feedback (RLHF), typically caps its responses to summarizations under 300 tokens, even in contexts of substantial length. This behavior is a contributing factor to GPT-4’s underachievement in BookSum, where more extensive and detailed responses are crucial for capturing the full scope of the content.
@@ 5.7 Ablation Study
We report the effect of Atomic Fact Chain Generation and Document-level Relation Extraction on the number of atomic facts to be verified.
@@
Table 5: Number of facts(or fact chains) to be evaluated under three settings, averaging 9 models we used for experiments, i.e. Atomic Fact Generation(AFG), Atomic Fact Chain Generation(AFCG) and Atomic Fact Chain Generation Complemented with DocRE(AFCG + DocRE).
From Table 5, datasets with longer context and richer information tend to result in more facts, and by using Atomic Fact Chain Generation, we reduce the number of cascaded facts a lot. By doing so, minor mistakes in one piece of information will not affect others. Also, DocRE complements AFG with document-level relations, which is helpful for the evaluation of the hierarchical information in the summary.
5.8 Meta Evaluation of Human Annotation
By observing the results of human annotations, we find that ground truth prefers to summarize more elegantly, using core arguments from the original context in addition to examples supported with other a little bit trivial arguments. However, LLMs try to list all of the arguments, ignoring the examples or data in the context, leading to differences with ground truth. However, there does exist some verbosity in the ground truth, indicating the natural drawback of the summarization dataset.
@@ 5.9 Analysis of Different LLM Verifier
Since our LLM Verifier should consider previous states in a logic chain and verify whether the current atomic fact is true or not, it is a rather hard problem compared with simply verifying. Intuitively, with more capable models, we will get a more human-aligned evaluation metric. We report the results using ChatGPT for a simple proof.
@@
EVA-Score-Mistral-7B-Instruct 0.710 0.673 0.503
@@
Table 6: Human correlation using ChatGPT as EVA-Score LLM verifier compared with EVA-Score with Mistral-7B-Instruct. The larger one is denoted is bold.
Leveraging ChatGPT results in more human correlation in Pearson and Kendall correlation and comparable Spearman correlation, which is consistent with our recognition. However, using ChatGPT results in much more costs than open-source LLMs because of the great number of validations.
@@ 5.10 Case Study
To provide a more intuitive and straightforward insight into the pipeline of EVA-Score, we provide a detailed case in Appendix B. From the case, we can see that EVA-Score provides not only a quantitative and objective information score but also provides a human-like explanation.
@@ 6 Conclusion and Future Work
In this work, we propose a new evaluation metric called EVA-Score to evaluate informativeness in long-form summarization. It extracts atomic facts which are then reformed as logic chains and document-level relations from reference and candidate and uses LLMs for validation one by one. Experiment results show that compared with other traditional metrics or LLM-based metrics, our work shows the best human correlation as well as provides an objective explanation. Moreover, we re-evaluate long-form summarization based on EVA-Score, indicating a great challenge in book-level summarization.
We believe that EVA-Score is meant to calculate information overlap between reference and candidate, thus it can be applied to broader scenarios rather than long-form summarization only. In future work, we suggest: 1) Leveraging EVA-Score to evaluate more reference-based and information-rich tasks and 2) Utilizing EVA-Score for the generation of more informative texts.
@@
EVA-Score can be time-consuming to run
EVA-Score includes several pre-processing methods. We should first extract atomic facts and document-level relations two summaries and use LLMs for paraphrasing to make the relations more natural. The pre-processing takes around two minutes per record. Then for each atomic fact in the candidate, we select $k$ most relevant from the reference and use LLMs for validation. Since the number of atomic facts takes quite a lot of time. We report the time consumed correlated to the sum of the length of the candidate and reference.
Length Interval/token 0-300 300-600 600-900 900-1200
@@
Table 7: Time consumed regarding to different summary length
@@
Some existing evaluation metrics consider many other factors, like faithfulness, and consistency. EVA-Score may be combined with other evaluation metrics to provide a more comprehensive evaluation of long-form summarization.
@@
EVA-Score implements Document-level Relation Extraction for the lack of hierarchical information. However, Doc-RE by GPT-4 maybe incomplete because GPT-4 may fail to capture the core meaning of the document level due to the difficulty of such a task. Leveraging more capable models that are fine-tuned on Doc-RE may be more helpful.

### v3

@@ 5 Experiment
In this section, we detail our implementations and the results of the experiments.
@@ 5.1 Dataset
To evaluate the applicability of EVA-Score, we gather long-form summarization across varying context lengths and domains. We adopt CNN Dailymail (See et al., 2017), PubMed, GovReport (Cohan et al., 2018), arXiv, and BookSum (Kryściński et al., 2022). In particular, we select the chapter-level summarization version of BookSum for less difficulty. For all datasets, we use validation datasets for evaluation. We list a detailed description of the dataset in Appendix B.1.
@@ 5.2 Baseline
We compare the results of EVA-Score with traditional evaluation metrics, i.e., ROUGE-1, ROUGE-2 (Lin, 2004) and BERTScore (Zhang et al., 2020) which are based on text overlap or word embeddings. For LLM-based evaluation metrics, we include ChatGPT and GPT-4-0125-preview as our evaluators since they are always considered one of the strongest LLMs, and we denote these methods as prompt-only metrics. Moreover, we validate on fine-tuned LLMs including Auto-J-13B (Li et al., 2023) and Critique-LLM-6B (Ke et al., 2024).We denote these two metrics as fine-tuned metrics. For ChatGPT and GPT-4, we instruct them using the prompt from Table 17. For all the LLMs, we set temperatures at $1.0$.
@@ 5.3 Human Annotation
We aim to develop an evaluation metric that resonates with human judgment. To achieve this, we randomly select 50 records from the datasets mentioned in Section 5.1 and assess the informativeness of each record. We employ five annotators, including some of the authors, all of whom are either M.S. or Ph.D. students. Initially, the annotators identify atomic facts at the sentence and document level, denoted as $DocS$ and $\hat{DocS}$, from $Sum$ and $\hat{Sum}$. They then determine the overlap between $DocS$ and $\hat{DocS}$. Redundant information, which appears multiple times, is recorded only once. Subsequently, they calculate Precision and Recall for these overlaps and derive the F1 Score as their harmonic mean. Furthermore, we ensure that each record is evaluated by at least two annotators to maintain the fairness of the assessment. Each annotation costs about $15$ minutes. The high time cost implies the complexity and high quality of our annotation. The detailed annotation recipe is listed in Appendix B.3. After the annotation process, we assessed the quality by measuring the agreement between the annotators. We achieve an Intraclass Correlation Coefficient (ICC) of 0.999 using a two-way random-effects model, and a Pearson correlation of 0.814. These high correlations demonstrate the robustness and reliability of our annotations.
@@ 5.4 Main Results
Trained Evaluation Metric
Auto-J-13B 0.285 0.233 0.167 0.885 0.947 0.913
CritiqueLLM-6B 0.510 0.490 0.353 0.865 0.800 0.667
Prompt-only Evaluation Metric
ChatGPT 0.135 0.210 0.162 0.430 0.600 0.333
@@
EVA-Score 0.710 0.673 0.503 0.975 1.000 1.000
Table 1: Text-level and system-level Pearson($\sigma$), Spearman($\gamma$), and Kendall($\tau$) correlations between different evaluation metrics and human annotations. The highest correlation among the methods is denoted in bold. Our EVA-Score achieves the best performance with a perfect correlation at the System-Level.
In this section, we report the correlation between each evaluation metric and human annotations. We adopt text-level and system-level Pearson ($\sigma$), Spearman ($\gamma$), and Kendall ($\tau$) correlation. Text-level correlation is calculated by computing the correlation coefficients between human evaluations and automatic evaluation metrics and then averaging these coefficients for all generated texts. System-level correlation, on the other hand, is determined by computing the correlation coefficients between human evaluations and automatic evaluation metrics for the average score of each dataset. The results are listed in Table 1. From the results, we observe that our EVA-Score outperforms all other traditional and LLM-based evaluation metrics in achieving the highest correlation with human judgments across all indicators at both levels. At the System Level, EVA-Score achieves perfect Spearman and Kendall correlations of 1.00, demonstrating complete alignment with human assessments. In contrast, Traditional Evaluation Methods, while somewhat correlated with human evaluations, generally lack strong correlations, highlighting their limitations. Models such as ChatGPT and GPT-4 exhibit a relatively low correlation with human annotations, which we attribute to the complexity involved in assessing the informativeness of a given summary against a reference, requiring them to decompose the task and finish the task step by step. Interestingly, evaluation models fine-tuned specifically for NLG achieve a high correlation with human annotations at the System Level. This improved performance is likely due to their fine-tuning process, which explicitly includes informativeness as an evaluation criterion, suggesting a promising direction for the future development of larger evaluation models.
5.5 Error Analysis
@@
Table 2: The results of several LLMs on different long-form summarization datasets. We denote EVA-Score as EVA, R-1 as ROUGE-1, and BS as BERTScore. All the values are normalized between 0 and 1. We indicate the highest score of EVA-Score using bold.
To have a thorough understanding of EVA-Score, we provide error analysis from the pipeline of EVA-Score in Appendix B.5 showing several situations where EVA-Score fails. To step further into why ROUGE and BERTScore are more suitable in relatively shorter settings than in longer settings, we provide our explanation in Appendix B.6.
5.6 Re-evaluation on Abstractive Long-form Summarization
To have a thorough understanding of the performance of LLMs on abstractive long-form summarization, we test on a wide range of models. For open source models, we evaluate on Mistral-7B-instruct (Jiang et al., 2023b), Llama-2-chat-7B, Llama-2-chat-13B (Touvron et al., 2023), Vicuna-7B, Vicuna-13B (Zheng et al., 2023), and Gemma-7B-it (Team et al., 2024). For close-source models, we experiment with ChatGPT and GPT-4-0125-preview. For all the LLMs, we set the temperatures as $1.0$. We use three different evaluation metrics to indicate their performance, i.e. EVA-Score, ROUGE-1, and BERTScore. Since this work focuses on the evaluation of long-form summarization, we do not pursue feeding long context into LLMs without information loss. Instead, we believe that useful information usually appears at the beginning or end of a text as Bai et al. (2023), and we simply truncate the context from the beginning and end to meet the maximum input sequence length of the models. The length of the context window can be viewed as part of the models’ inherent abilities, so it will not introduce unfairness in our evaluation. For all GLMs, our prompt is “Summary the following text. You can use your own words instead of extracting from the original document.”, ensuring an abstractive setting.
Referring to Table 2, as context length increases, all models experience varying degrees of performance decline, particularly in the BookSum dataset, where results are notably poorer. However, models capable of handling longer input sequences (Su et al., 2023) demonstrate superior performance in abstractive long-form summarization compared to those with limited context windows when dealing with larger contexts, largely due to their ability to comprehend extensive content. Conversely, when the context length is shorter, these models perform comparably or even worse than models with normal sequence lengths, such as Vicuna-7B. GPT-4 demonstrates robust capabilities compared to others, particularly in situations of longer input sequences under EVA-Score. However, it fails in relatively shorter settings, probably resulting from the tendency to respond in more words than others. Surprisingly, Mistral-7B-instruct, compared with other models with the same amount of parameters or even more parameters, shows great potential to fulfill abstractive long-form summarization and even beats GPT-4 by a small margin in BookSum, which meets the findings of Chang et al. (2024).
Moreover, compared to other evaluation metrics, EVA-Score shows a more pronounced and consistent decrease in performance as the context length grows, with fewer fluctuations. This pattern mirrors human cognitive processes, where increasing information load makes maintaining summarization quality more challenging. The consistent decline highlights the effectiveness of EVA-Score from another perspective.
@@ 5.7 Analysis of Summarization Length
Metric Avg Length Avg Length Avg Length Avg Length Avg Length
@@
Llama-2-7B-chat 118 143 108 140 185
Llama-2-13B-chat 124 184 221 236 179
@@
Table 3: The average length of summary counted in words generated by different models denoted by Avg Length. The largest count is marked using bold. It can be seen that GPT-4 prefers to summarize with the length of responses in a narrow range.
To elucidate why GPT-4 underperforms under EVA-Score in the situation of a shorter context, we analyze the summary lengths produced by different models, as detailed in Table 3.
We observe that shorter contexts typically result in shorter ground truth. Consequently, if a model generates a lengthy summary, this verbosity is penalized with a reduced EVA-Score due to decreased precision. As the context lengthens, the references become richer in information. If these models tend to maintain a similar response length, lacking significant information addition, they fail to capture enough relevant points. GPT-4, which is refined through RLHF (Ouyang et al., 2022), typically caps its responses to summarizations to a narrow range. This behavior contributes to GPT-4’s underachievement in Booksum, where more extensive and detailed responses are crucial to capture the full scope of the content. Additionally, we provide a meta-evaluation of human annotations in Appendix B.4, highlighting the underlying approach adopted by humans and models for summarization.
@@ 5.8 Ablation Study
We report the effect of AFCG and DocRE on the number of atomic facts to be verified.
@@
Table 4: Number of facts (fact chains) to be evaluated under three settings, averaging 9 models we used for experiments, i.e., Atomic Fact Generation (AFG), Atomic Fact Chain Generation (AFCG) and Atomic Fact Chain Generation Complemented with Document-level Relation Extraction (AFCG + DocRE).
From Table 4, datasets with longer context and richer information tend to result in more facts, and by AFCG, we largely reduce the number of cascaded facts. Also, DocRE complements AFG with document-level relations, which is helpful for the evaluation of the hierarchical information in the summary. The two processes work together to form a concise yet comprehensive list of facts to be verified.
@@ 5.9 Analysis of Different LLM Verifier
Since our LLM Verifier should consider previous states in a logic chain and verify the accuracy of current atomic fact, it is a rather hard problem compared to simple verification of an NLI problem. Intuitively, with stronger models, we will get a more human-aligned evaluation metric. Therefore, we report the results using ChatGPT for a simple proof which are listed in Table 5.
Leveraging ChatGPT results in a higher human correlation in Pearson and Kendall correlation and comparable Spearman correlation, which is consistent with our recognition. However, using ChatGPT results in much more costs than open-source LLMs because of the great number of validations.
@@
EVA-Score-Mixtral-7B-Instruct 0.710 0.673 0.503
@@
Table 5: Human correlation using ChatGPT as LLM verifier of EVA-Score compared with EVA-Score with Mistral-7B-Instruct. $\sigma$, $\gamma$, $\tau$ are Pearson, Spearman, and Kendall correlation respectively. The larger one is denoted in bold.
@@ 5.10 Case Study
To provide a more intuitive and straightforward insight into the pipeline of EVA-Score, we provide a detailed example in Appendix A.5. From the case, we can see that EVA-Score provides not only a quantitative and objective information score but also provides a human-like explanation.
@@ 6 Conclusion and Future Work
In this work, we introduce a novel evaluation metric, termed EVA-Score, to evaluate the informativeness in abstract long-form summarization. EVA-Score begins by extracting atomic facts from given summaries, which are then restructured into logical chains. We also extract document-level relations for the completion of information. Subsequently, LLMs are employed for further validation of the information. Experimental results demonstrate that, compared to existing automatic evaluation metrics, EVA-Score achieves the highest correlation with human judgments, providing an objective and explainable assessment. Furthermore, we use EVA-Score to re-evaluate abstractive long-form summarization models, highlighting the challenges LLMs face in handling extended contexts. We also conduct a detailed analysis of cases where EVA-Score and other evaluation metrics fall short in evaluating abstractive long-form summaries. We believe that EVA-Score, designed to quantify information overlap between reference and candidate, has potential applications beyond long-form summarization and can be extended to broader evaluation scenarios.
@@
EVA-Score is time-consuming to run
EVA-Score includes several pre-processing methods. We should first extract atomic facts and document-level relations from two summaries. We also use LLMs for paraphrasing to make the relations more natural. The preprocessing takes around two minutes per record. Then for each atomic fact in the candidate, we select $k$ most relevant from the reference and use LLMs for validation. Since the number of atomic facts is very large, it takes quite a lot of time. We report the time consumed correlated to the sum of the length of the candidate and reference counted in tokens.
Length Interval 0-300 300-600 600-900 900-1200
@@
Table 6: Time consumed measured in seconds for different summary lengths measured in tokens. The time consumption shows an approximately linear correlation with text length.
@@
EVA-Score is designed to evaluate informativeness. Some existing evaluation metrics consider many other factors, like faithfulness, and consistency. EVA-Score may be combined with other evaluation metrics to provide a more comprehensive and fair evaluation of abstractive long-form summarization.
@@
EVA-Score incorporates DocRE to address the lack of hierarchical information. However, DocRE by GPT-4 may be incomplete, as GPT-4 may struggle to fully capture the core meaning at the document level due to the inherent complexity of the task. Leveraging more advanced models specifically fine-tuned for DocRE could potentially improve results. Furthermore, the filter used in the current approach is simplistic, and experimental findings suggest that a more effective filter is needed to better distinguish document-level relations from sentence-level ones.